Assume that human body temperatures are normally distributed with a mean of 98.20 degrees Upper F and a standard deviation of 0.61 degrees Upper F.

a. A hospital uses 100.6 degrees Upper F as the lowest temperature considered to be a fever. What percentage of normal and healthy persons would be considered to have a​ fever? Does this percentage suggest that a cutoff of 100.6 degrees Upper F is​ appropriate?

b. Physicians want to select a minimum temperature for requiring further medical tests. What should that temperature​ be, if we want only​ 5.0% of healthy people to exceed​ it? (Such a result is a false​ positive, meaning that the test result is​ positive, but the subject is not really​ sick.) Click to view page 1 of the table.LOADING... Click to view page 2 of the table.LOADING... a. The percentage of normal and healthy persons considered to have a fever is
Answer #1

a) P(X > 100.6)

= P(z > (100.6 - 98.20)/0.61)

= P(z > 3.93)

= 0.0000

b) z score for top 5% area = 1.645

Hence,

Required temperature = 98.20 + 1.645*0.61 = 99.20